the length of each leg of an isosceles triangle is 5 centemeters more than twice the length of the base. if the perimeter of this isosceles triangle is 95 centemeters, what is the length of the base?.

Answer by stanbon(75887)

the length of each leg of an isosceles triangle is 5 centemeters more than twice the length of the base. if the perimeter of this isosceles triangle is 95 centemeters, what is the length of the base?.
--
Let base length be "x".
Each leg is "2x+5"
---
Equation:
x + 2(2x+5) = 95
x + 4x + 10 = 95
5x = 85
x = 17 cm (base length)
2x+5 = 39 cm (length of each base)
=======================
Cheers,
Stan H
================